# d3LLM: Ultra-Fast dLLM using Pseudo-Trajectory Distillation

### v1

@@ 1 Introduction
Diffusion large language models (dLLMs) have emerged as a promising alternative to autoregressive (AR) LLMs. A key advantage of dLLMs is their use of bidirectional attention, which enables capabilities such as parallel decoding, error correction, and random-order generation—features that are not feasible with AR models. Recently, several closed-source diffusion models, including Mercury (Labs et al., 2025), Gemini Diffusion (Google DeepMind, 2025), and Seed Diffusion (Song et al., 2025), have demonstrated impressive efficiency and performance, achieving extremely high throughput and sometimes exceeding 1000 tokens per second in certain settings. In contrast, open-source dLLMs have exhibited significantly lower throughput, sometimes even slower than AR baselines. For example, LLaDA (Nie et al., 2025) and Dream (Ye et al., 2025) achieve only around 20 tokens per second. Moreover, they often lag behind similarly-sized AR models in terms of accuracy.
@@ 2 Problem Formulation
The intuition behind AUP is simple: (i) If you increase parallelism without losing accuracy, your AUP increases a lot. (ii) If you increase parallelism by sacrificing accuracy, your AUP increases only a little (or not at all), because the penalty suppresses low-accuracy regimes. AUP thus provides a unified measure of decoding quality under increasing parallelism, encouraging models to achieve a fair balance between accuracy and parallelism.
@@
The hyperparameter $\alpha$ controls the penalty for accuracy degradation. A larger $\alpha$ increases sensitivity to performance drops, causing the contribution of throughput to decay exponentially with the error rate. In the ideal case, where a method improves parallelism without compromising accuracy, the AUP reduces to the standard area under the parallelism-accuracy curve (AUC). In our setting, we set $\alpha=3$ as it balances the importance of parallelism and accuracy.
@@
AUP is hardware-independent because it is built on TPF rather than TPS. TPS is heavily affected by hardware generation (H100 vs. A100), kernel fusion, cache implementation, and the inference framework. The same algorithm can appear dramatically different depending on system details. For instance, in our experiments, our d3LLM-LLaDA model (introduced in Section 3) achieves approximately $5\times$ higher TPS than an AR baseline (Qwen-2.5-7B-it) on an NVIDIA H100 GPU. However, this advantage shrinks significantly on an NVIDIA A100 GPU. In contrast, TPF captures algorithmic parallelism: how many tokens are generated per forward pass, which is much more stable across hardware. Therefore, AUP gives a fairer view of algorithmic progress, without requiring everyone to run on the exact same GPU or inference engine, helping the community focus on algorithmic design without requiring access to particular GPUs.
@@
As discussed in the introduction, existing dLLMs face two key limitations: (i) standard training with random masking lacks guidance on which tokens can be decoded earlier, resulting in suboptimal unmasking orders; (ii) while parallel decoding improves parallelism, it often degrades generation quality due to error propagation from incomplete context. To address these limitations and further push the accuracy-parallelism frontier, we propose d3LLM (pseuDo-Distilled-Diffusion Large Language Model), a framework that improves both training and inference:
@@
    Pseudo-trajectory distillation (training): We distill from teacher dLLM’s decoding trajectory, providing intermediate supervision on token unmasking order, combined with curriculum learning that gradually increases difficulty.
@@
    Multi-block decoding with KV-refresh (inference): We parallel decode multiple blocks based on entropy, with a KV-refresh mechanism to mitigate quality degradation.
In the following, we detail each component.
@@ 3.1 Pseudo-Trajectory-Based Distillation Recipe
We propose a novel pseudo-trajectory-based distillation method to accelerate diffusion language models (dLLMs) through more informed training. This approach introduces an advanced distillation recipe aimed at improving both decoding efficiency and alignment with the teacher model’s generation strategy. Specifically, it incorporates the following key technique:
@@
Specifically, given a prompt $\mathbf{x}$ and a predefined maximum output length $n$, we first let the teacher dLLM generate and record its own decoding trajectory $\{\mathcal{T}_{1},\ldots,\mathcal{T}_{n}\}$, where $\mathcal{T}_{i}\in\mathbb{R}^{n},\ \forall i\in\{1,\ldots,n\}$. Rather than relying on the content of the teacher’s response, we extract only the order in which tokens are decoded. This order forms what we refer to as the pseudo-trajectory of the teacher. We combine the pseudo-trajectory with the ground-truth prompt–response pair $(\mathbf{x},\mathbf{y})$ and construct a noisy sequence $\widetilde{\mathbf{y}}\in\mathbb{R}^{n}$ that simulates the teacher’s intermediate state during the decoding process. Formally, let $t\in[0,1]$ denote the mask ratio, and $w=\{s,\ldots,s+k\}$ be a decoding window of length $k$. The noisy sequence $\widetilde{\mathbf{y}}$ is
@@ 3.2 Multi-Block Decoding Strategy
In addition to the novel distillation recipe, we also introduce a more efficient decoding mechanism tailored in d3LLM, designed to maximize parallelism across multiple decoding blocks. Our decoding recipe includes:
@@
Each block can be in one of the five states: Inactive, Activated, Fully-Activated, Completed but Stabilizing, and Completed. Transition rules are as follows: we create a new Activated block when its preceding block reaches $10\%$ completion and employ a conservative decoding strategy for this block, generating tokens only when they meet the entropy threshold. When the preceding block reaches $95\%$ completion, the Activated block transitions to a Fully-Activated state, where a more aggressive strategy is used by decoding at least one token per forward pass, regardless of the threshold. Once all tokens in a block are unmasked, the block enters the Completed but Stabilizing state, during which we perform forward passes without using the KV cache and refresh previous caches. After 1 or 2 rounds, the block becomes Completed, and we store its KV cache. In addition, we apply a periodic-refresh strategy that updates the KV cache every few rounds. This multi-block decoding strategy increases TPF by 30%, and the KV-refresh mechanism helps maintain the accuracy.
@@ 4 Experiments
  • Q1. Is the metric AUP reasonable?
  • Q2. How effective is d3LLM in terms of AUP score?
  • Q3. Is each module in d3LLM effective?
@@ 4.1 Experimental Setup
Contenders.   To validate the effectiveness of our approach, we compare d3LLM framework with state-of-the-art dLLM methods, including vanilla LLaDA (Nie et al., 2025) and Dream (Ye et al., 2025), training-free inference acceleration method Fast-dLLM (Wu et al., 2025b), causal block-diffusion methods Fast-dLLM-v2 (Wu et al., 2025a), dParallel (Chen et al., 2025), and a dLLM distillation method D2F (Wang et al., 2025a). More details about the contenders are deferred in Appendix A.4.
@@
Benchmark Datasets.   We present benchmark results across five representative tasks: GSM8K-CoT (chain-of-thought reasoning) (Gao et al., 2024), MATH (mathematical problem solving) (Lewkowycz et al., 2022), HumanEval (code generation) (Chen et al., 2021), MBPP (Python programming) (Austin et al., 2021b), and a long-context math reasoning task (5-shot GSM8K (Cobbe et al., 2021), with a prompt length $\approx$ 1000). These datasets span diverse domains and problem types and are widely used in the research community. In addition, their relatively long output lengths allow us to effectively evaluate the models’ parallel decoding capabilities together with their accuracy. These evaluations assess the performance of our proposed model, d3LLM, against state-of-the-art diffusion-based language models using three key metrics: parallelism, measured by tokens per forward pass (TPF), accuracy (solve rate / pass@1 accuracy depending on the benchmark), and our proposed AUP (Accuracy Under Parallelism) score.
@@ 4.2 Evaluation of d3LLM Framework
Results on LLaDA-based Models.   As shown in Table 1, our d3LLM-LLaDA consistently achieves the highest AUP scores across all five benchmark tasks, demonstrating the effectiveness of our proposed framework. Specifically, d3LLM-LLaDA achieves an AUP score of 635.7 on GSM8K-CoT, significantly outperforming the second-best method dParallel (358.1). On MATH dataset, d3LLM-LLaDA achieves 111.7 AUP, which is 73.2% higher than dParallel (64.5). Similar improvements are observed on HumanEval and Long-GSM8K. This superior performance can be attributed to d3LLM’s pseudo-trajectory distillation, which enables the model to learn the teacher’s token unmasking order, and the multi-block decoding strategy, which maximizes parallelism while maintaining accuracy through the KV-cache refresh mechanism.
@@
Results on Dream-based Models.   As shown in Table 2, d3LLM-Dream achieves the highest AUP scores on 4 out of 5 tasks, further validating the generalizability of our framework. For example, on GSM8K-CoT, d3LLM-Dream achieves an AUP of 400.4, outperforming dParallel (245.7) and Fast-dLLM-v2 (156.0). On MBPP-Instruct, d3LLM-Dream achieves an AUP score of 134.7, which is 24.7% higher than dParallel (108.0). Similar improvements are observed on other datasets, and the radar chart in Figure 4(c) further demonstrates that d3LLM-Dream achieves the largest overall coverage area among Dream-based methods, indicating its balanced superiority across diverse benchmarks. These results affirmatively answer Q2: our d3LLM framework is effective across different base models and tasks.
@@
To summarize, our d3LLM framework achieves the highest AUP scores with negligible performance degradation, successfully striking a balance between accuracy and parallelism. It delivers up to $5\times$ speedup over autoregressive decoding (Qwen-2.5-7B-it) on H100 GPUs and approximately $3.6\times$ speedup on A100 GPUs. Due to page limit, we defer more experimental results of about our efficient dLLM-coder, d3LLM-Coder, to Appendix A.
@@ 4.3 Ablation Study
Ablation Study on Distillation Recipe.   To validate each component’s contribution in the distillation recipe, we conduct ablation studies on d3LLM-LLaDA and evaluate it on the GSM8K-CoT dataset. As shown in Table 5 (upper part), we compare our full method with three variants: (i) a baseline model with only multi-block decoding and early stopping but without any distillation components, (ii) a model with pseudo-trajectory distillation only, and (iii) a model with pseudo-trajectory and curriculum noise but without curriculum windows. The baseline without distillation achieves a TPF of 6.41 with 72.2% accuracy. Adding pseudo-trajectory distillation improves TPF by 17.8% (from 6.41 to 7.55) while maintaining similar accuracy, demonstrating that learning the teacher’s token unmasking order effectively improves parallelism. Incorporating curriculum noise further increases TPF to 8.46 (12.1% improvement), though with an accuracy drop to 69.8%, indicating that the curriculum learning strategy enables more aggressive parallel decoding. Finally, adding curriculum window size yields our full model with TPF of 9.11 and accuracy of 73.1%, achieving a 7.7% TPF improvement while recovering accuracy. This demonstrates that the curriculum window strategy not only improves parallelism but also stabilizes the distillation process.
Ablation Study on Decoding Strategy.   As shown in Table 5 (lower part), we further ablate the decoding strategy components on d3LLM-LLaDA. Starting from our full distillation recipe, we compare: (i) vanilla block diffusion decoding without multi-block or early stopping, (ii) multi-block decoding without early stopping, and (iii) our complete decoding strategy with both multi-block decoding and early stopping. The multi-block decoding strategy enables parallel decoding across multiple blocks based on token entropy, which significantly improves TPF by allowing confident tokens in future blocks to be decoded simultaneously with the current block. The early stopping mechanism further optimizes efficiency by terminating decoding upon generating the EOS token, eliminating redundant forward passes. Together, these components achieve a TPF of 9.11 with 73.1% accuracy, yielding the highest AUP score of 635.7.
@@
Hyperparameter Analysis.   We further investigate the impact of key hyperparameters in our curriculum learning strategy: the curriculum noise level and the curriculum window size strategy. As shown in Table 7, using a fixed noise level ($t=0.5$) achieves a TPF of 7.49 and accuracy of 72.8%, while our curriculum noise strategy (0.0 $\rightarrow$ 0.8) improves TPF to 9.11 with accuracy of 73.1%, yielding a 21.6% improvement in TPF and 22.2% improvement in AUP score. This validates that gradually increasing the noise level during training enables the model to first learn basic token dependencies before handling more challenging masking patterns. Similarly, Table 7 shows that a fixed window size (kk=32) achieves a TPF of 8.22 with accuracy of 69.8%, while our curriculum window strategy (16 $\rightarrow$ 32) improves TPF to 9.11 with accuracy of 73.1%, yielding a 19.0% improvement in AUP score. Interestingly, starting from too small a window (0 $\rightarrow$ 32) leads to lower accuracy (72.8%) and AUP score (603.1), suggesting that an appropriate initial window size is crucial for stable distillation. These results demonstrate that our curriculum learning strategy with properly tuned schedules effectively balances parallelism and accuracy.

### v2

# d3LLM: Ultra-Fast Diffusion LLM using Pseudo-Trajectory Distillation

@@ Abstract
Diffusion Large Language Models, Fast Inference, Parallel Decoding, Efficiency
[…]
@@ 1 Introduction
Diffusion large language models (dLLMs) have emerged as a promising alternative to autoregressive (AR) LLMs. A key advantage of dLLMs is their use of bidirectional attention, which enables capabilities such as parallel decoding, error correction, and random-order generation—features that are not feasible with AR models. Recently, several closed-source diffusion models, including Mercury (Khanna et al., 2025), Gemini Diffusion (Google DeepMind, 2025), and Seed Diffusion (Song et al., 2025), have demonstrated impressive efficiency and performance, achieving extremely high throughput and sometimes exceeding 1000 tokens per second in certain settings. In contrast, open-source dLLMs have exhibited significantly lower throughput, sometimes even slower than AR baselines. For example, LLaDA (Nie et al., 2025) and Dream (Ye et al., 2025) achieve only around 20 tokens per second. Moreover, they often lag behind similarly-sized AR models in terms of accuracy.
@@ 2 Problem Formulation
The intuition behind AUP is simple: (i) If parallelism is increased without losing accuracy, AUP increases a lot. (ii) Instead, if parallelism is increased by sacrificing accuracy, AUP increases only a little, as the penalty suppresses low-accuracy regimes. AUP thus provides a unified measure of quality under increasing parallelism, encouraging to achieve a fair balance between accuracy and parallelism.
@@
The hyperparameter $\alpha$ controls the penalty for accuracy degradation. A larger $\alpha$ increases sensitivity to performance drops, causing the contribution of parallelism gains to decay exponentially with increasing error rate. In the ideal case where a method improves parallelism without compromising accuracy, AUP reduces to the standard area under the parallelism-accuracy curve (AUC). We set $\alpha=3$ as the default, as it provides a reasonable balance between parallelism and accuracy. We validate different choices of $\alpha$ in Appendix A.4.
@@
AUP is hardware-independent because it is built on TPF rather than TPS. TPS is heavily affected by hardware generation, kernel fusion, and the inference framework. The same algorithm can appear dramatically different depending on system details. For instance, in our experiments, our d3LLM-LLaDA model (introduced in Section 3) achieves approximately $5\times$ higher TPS than an AR baseline (Qwen-2.5-7B-it) on an NVIDIA H100 GPU. However, this advantage shrinks significantly on an NVIDIA A100 GPU. In contrast, TPF captures algorithmic parallelism: how many tokens are generated per forward pass, which is much more stable across hardware. Therefore, AUP gives a fairer view of algorithmic progress, without requiring everyone to run on the exact same GPU or inference engine, helping the community focus on algorithmic design without requiring access to particular GPUs.
@@
This section introduces our approach, d3LLM framework.
Motivation.   As demonstrated in the previous section, dLLMs inherently face an accuracy–parallelism trade-off. While AUP cannot be directly optimized as a training objective (see Appendix A.4), it provides a guiding principle: balancing accuracy and parallelism is essential. Based on this insight, we propose d3LLM (pseuDo-Distilled-Diffusion Large Language Model), a framework that improves both training and inference recipes:
@@
    Pseudo-trajectory distillation (training): We incorporate the teacher dLLM’s own decoding trajectory into the distillation process to provide intermediate supervision indicating which tokens can be safely decoded earlier, thereby improving parallelism. We further employ a curriculum learning strategy that progressively increases difficulty, enabling stable training and preserving accuracy.
@@
    Multi-block decoding with KV-refresh (inference): We introduce entropy-based multi-block decoding that simultaneously decodes multiple blocks by prioritizing high-confidence tokens, thereby improving parallelism. To mitigate quality degradation, we employ a KV-cache refresh mechanism that periodically recomputes all previously cached states, thereby preserving accuracy.
@@ 3.1 Pseudo-Trajectory-Based Distillation Recipe
We propose a novel pseudo-trajectory-based distillation recipe. Specifically, it incorporates following key technique:
@@
Specifically, given a prompt $\mathbf{x}$ and a predefined maximum output length $n$, we first use the teacher dLLM to generate and record its own decoding trajectory $\{\mathcal{T}_{1},\ldots,\mathcal{T}_{n}\}$, where $\mathcal{T}_{i}\in\mathbb{R}^{n},\ \forall i\in\{1,\ldots,n\}$. Here, we constrain the teacher model to unmask exactly one token at each decoding step, and we continue generation beyond the EOS token so that the output length is exactly $n$. Rather than relying on the content of the teacher’s response, we extract only the order in which tokens are decoded, which we refer to as the pseudo-trajectory of the teacher. We combine it with the ground-truth prompt–response pair $(\mathbf{x},\mathbf{y})$ and construct a noisy sequence $\widetilde{\mathbf{y}}\in\mathbb{R}^{n}$ that simulates teacher’s intermediate state during the decoding. Formally, let $t\in[0,1]$ denote the mask ratio, and $w=\{s,\ldots,s+k\}$ be a decoding window of length $k$, the noisy sequence $\widetilde{\mathbf{y}}\in\mathbb{R}^{n}$ is
@@ 3.2 Multi-Block Decoding Strategy
In addition to the novel distillation recipe, we also introduce a parallel decoding strategy, designed to maximize parallelism across multiple blocks, and preserving the accuracy.
@@
Each block can be in one of the five states: Inactive, Activated, Fully-Activated, Completed but Stabilizing, and Completed. Transition rules are as follows: we create a new Activated block when its preceding block reaches $10\%$ completion and employ a conservative decoding strategy for this block, generating tokens only when they meet the entropy threshold. When the preceding block reaches $95\%$ completion, the Activated block transitions to a Fully-Activated state, where a more aggressive strategy is used by decoding at least one token per forward pass, regardless of the threshold. Once all tokens in a block are unmasked, the block enters the Completed but Stabilizing state, during which we perform forward passes without using the KV cache and refresh previous caches. After 1 or 2 rounds, the block becomes Completed, and we store its KV cache. In addition, we apply a periodic-refresh strategy that updates the KV cache every few rounds. This multi-block decoding strategy increases TPF by 30%, and the KV-cache refresh mechanism helps maintain the accuracy.
@@ 4 Experiments
  • Q1.
    Is the metric AUP reasonable?
  • Q2.
    How effective is d3LLM in terms of AUP score?
  • Q3.
    Is each module in d3LLM effective?
@@ 4.1 Experimental Setup
Contenders.   To validate the effectiveness of our approach, we compare d3LLM framework with state-of-the-art dLLM methods, including vanilla LLaDA (Nie et al., 2025) and Dream (Ye et al., 2025), training-free inference acceleration method Fast-dLLM (Wu et al., 2025b), causal block-diffusion methods Fast-dLLM-v2 (Wu et al., 2025a), dParallel (Chen et al., 2025), and D2F (Wang et al., 2025a).
@@
Benchmark Datasets.   We present benchmark results across five representative tasks: GSM8K-CoT (chain-of-thought reasoning) (Gao et al., 2024), MATH (mathematical problem solving) (Lewkowycz et al., 2022), HumanEval (code generation) (Chen et al., 2021), MBPP (Python programming) (Austin et al., 2021b), and a long-context math reasoning task (5-shot GSM8K (Cobbe et al., 2021), with a prompt length $\approx$ 1000). These datasets span diverse domains and problem types and are widely used in the research community. In addition, their relatively long output lengths allow us to effectively evaluate the models’ parallel decoding capabilities together with their accuracy. We assess the performance of different dLLM methods using three key metrics: parallelism, measured by the tokens per forward pass (TPF); accuracy (the solve rate / pass@1 accuracy depending on the benchmark); and our proposed AUP (Accuracy Under Parallelism) score.
@@ 4.2 Evaluation of d3LLM Framework
Results on LLaDA-based Models.   As shown in Table 1, our d3LLM-LLaDA consistently achieves the highest AUP scores across all five benchmark tasks, demonstrating the effectiveness of our proposed framework. Specifically, d3LLM-LLaDA achieves an AUP score of 637.7 on GSM8K-CoT, significantly outperforming the second-best method dParallel (358.1). On MATH dataset, d3LLM-LLaDA achieves 107.6 AUP, which is higher than dParallel (64.5). Similar improvements are observed on HumanEval and Long-GSM8K. This superior performance can be attributed to d3LLM’s pseudo-trajectory distillation, which enables the model to learn the teacher’s token unmasking order, and the multi-block decoding strategy, which maximizes parallelism while maintaining accuracy through the KV-cache refresh mechanism.
@@
Results on Dream-based Models.   As shown in Table 2, d3LLM-Dream achieves the highest AUP scores on 4 out of 5 tasks, further validating the generalizability of our framework. For example, on GSM8K-CoT, d3LLM-Dream achieves an AUP of 391.3, outperforming dParallel (245.7) and Fast-dLLM-v2 (156.0). On MBPP-Instruct, d3LLM-Dream achieves an AUP score of 141.4, which is higher than dParallel (108.0). Similar improvements are observed on other datasets, and the radar chart in Figure 4(c) further demonstrates that d3LLM-Dream achieves the largest overall coverage area among Dream-based methods, indicating its balanced superiority across diverse benchmarks. These results affirmatively answer Q2: our d3LLM framework is effective across different base models and tasks.
@@
To summarize, d3LLM achieves the highest AUP scores with negligible performance degradation, successfully striking a balance between accuracy and parallelism. It delivers up to $5\times$ speedup over autoregressive decoding (Qwen-2.5-7B-it) on H100 GPUs and approximately $3.6\times$ speedup on A100 GPUs. Due to page limit, we defer more experimental results of about d3LLM-Coder to Appendix A.
@@ 4.3 Ablation Study
Ablation Study on Distillation Recipe.   To validate each component’s contribution in the distillation recipe, we conduct ablation studies on d3LLM-LLaDA and evaluate it on the GSM8K-CoT dataset. As shown in Table 5 (upper part), we compare our full method with three variants: (i) a baseline model with only multi-block decoding and early stopping but without any distillation components, (ii) a model with pseudo-trajectory distillation only, and (iii) only without curriculum windows. The baseline without distillation achieves a TPF of 6.41 with 72.2% accuracy. Adding pseudo-trajectory distillation improves TPF by 17.8% (from 6.41 to 7.55) while maintaining similar accuracy, demonstrating that learning the teacher’s token unmasking order effectively improves parallelism. Incorporating curriculum noise further increases TPF to 8.46, though with an accuracy drop to 69.8%, indicating that the curriculum learning strategy enables more aggressive parallel decoding. Finally, adding curriculum window size yields our full model with TPF of 9.11 and accuracy of 73.1%, achieving a TPF improvement while recovering accuracy. This demonstrates that the curriculum window strategy not only improves parallelism but also stabilizes the distillation process.
Ablation Study on Decoding Strategy.   As shown in Table 5 (lower part), we further ablate the decoding strategy components on d3LLM-LLaDA. Starting from our full distillation recipe, we compare: (i) vanilla block diffusion decoding without multi-block or early stopping, (ii) multi-block decoding without early stopping, and (iii) our complete decoding strategy with both multi-block decoding and early stopping. The multi-block decoding strategy enables parallel decoding across multiple blocks based on token entropy, which significantly improves TPF by allowing confident tokens in future blocks to be decoded simultaneously with the current block. The early stopping mechanism further optimizes efficiency by terminating decoding upon generating the EOS token, eliminating redundant forward passes.
@@
Hyperparameter Analysis.   We further investigate the impact of key hyperparameters in our curriculum learning strategy: the curriculum noise level and the curriculum window size strategy. As shown in Table 7, using a fixed noise level ($t=0.5$) achieves a TPF of 7.49 and accuracy of 72.8%, while our curriculum noise strategy (0.0 $\rightarrow$ 0.8) improves TPF to 9.11 with accuracy of 73.1%, yielding a 21.6% improvement in TPF and 22.2% improvement in AUP score. This validates that gradually increasing the noise level during training enables the model to first learn basic token dependencies before handling more challenging masking patterns. Similarly, Table 7 shows that a fixed window size (kk=32) achieves a TPF of 8.22 with accuracy of 69.8%, while our curriculum window strategy (16 $\rightarrow$ 32) improves TPF to 9.11 with accuracy of 73.1%, yielding a 19.0% improvement in AUP score. Interestingly, starting from too small a window (0 $\rightarrow$ 32) leads to lower accuracy and AUP score, suggesting that an appropriate window size is crucial for stable distillation. These results demonstrate that our curriculum learning strategy with properly tuned schedules effectively balances parallelism and accuracy.